# PET: Preference Evolution Tracking with LLM-Generated Explainable Distribution

### v2

@@ 3.1 Problem Formulation
To ensure reproducibility and simplicity, we use this transparent and normalized frequency of a single action instead of heuristic weighting of different interaction types, as our framework is designed to be agnostic to the specific construction of the ground-truth label. Formally, we aim to minimize the expected divergence between these two distributions:
$\min_{\mathcal{M}}\mathbb{E}_{H_{1:t}}\left[\mathcal{D}\left(\bar{\theta}_{t+1:T}\parallel\hat{\theta}_{t+1:T}\right)\right],$ (4)
where $\mathcal{D}$ is a measure of distributional similarity, such as the Jensen-Shannon Divergence (JS-Div). The ranked list of preferences used for evaluation with metrics like NDCG is then obtained by sorting the elements of $\hat{\theta}_{t+1:T}$ in descending order.
3.2 Methodology
This section outlines the PET framework for learning user preferences, which consists of two core stages: model alignment and preference inference.
The alignment stage follows a structured two-phase training approach. First, the model undergoes large-scale pre-training (PT) on domain-specific corpora to establish foundational knowledge and capture population-level patterns. Subsequently, it is refined through supervised fine-tuning (SFT) on carefully curated user interaction sequences, effectively aligning its internal representations with the nuances of personalized preference prediction. The SFT data are constructed from each user’s full history by generating pairs (context, label), defined as follows: the context is an expanding historical window $H_{1:t}$, and the label is the user’s future preference distribution $\bar{\theta}_{t+1:T}$. This aligned model forms the basis for all subsequent preference inference methods.
We now present two inference methods for preference distribution estimation, which differ primarily in their querying strategy (more detailed design of the prompts and algorithms presented in Appendices A and B).
Method 1: Likelihood-based Probing. This approach decomposes a multi-class problem into a series of independent binary classifications. The LLM is used as a “prober” or “scorer.” For each cluster $c\in C$, we iteratively ask the LLM for inferences and extract the logit score of positive tokens and then apply a softmax function on the score vector to recover the inferred probability distribution (petroni2019language; schick2020s).
Method 2: Generative Classification. This method uses a single forward pass. A prompt is constructed that frames the task as a multi-choice question, with each cluster mapped to a unique token (e.g., ‘A’, ‘B’). We then read the model’s next-token logits for these specific cluster tokens and normalize them via a softmax function to obtain the final probability distribution (raffel2020exploring; chung2024scaling).
Extreme multi-label preference inference.
To scale PET to real-world applications involving thousands of fine-grained preference clusters – such as in large video platforms or e-commerce systems – we introduce a two-stage Hierarchical Probing strategy. Existing methods face clear limitations in such settings: Likelihood-based Probing exhibits linear computational complexity in the number of clusters $K$, becoming prohibitively expensive, while Generative Classification often yields unstable and semantically ungrounded logits when forced to assign probabilities to artificial class tokens.
To overcome these issues, we organize clusters into a two-level semantic taxonomy: coarse-grained L1 categories (e.g., Food & Drink, Technology) partition the full L2 label space into interpretable subgroups. This tree‑structured approach is also validated in recent extreme multi‑label classification work in other tasks (chen2023token; wan2023explore; zhou2024quest). Inference proceeds in two stages: (1) L1 Preference Scoping: estimates marginal probabilities over top-level categories; (2) Strategic L2 Exploration: selectively probes child clusters, guided by downstream goals. For example, one may prioritize top-ranked L1 branches for relevance, or intentionally explore long-tail branches for novelty. This decomposition improves both tractability and semantic alignment by guiding the LLM from general categories to relevant fine-grained clusters, enabling scalable and adaptive preference modeling. See the detailed computational complexity analysis in Appendix C.
3.3 Theoretical Foundations and Guarantees
This section provides a theoretical argument that PET’s methodology is optimal for standard ranking metrics compared to the direct generation paradigm. We begin with a key assumption about our model:
Assumption 1 (Isotonic Probing Estimator).
We assume that the vector of logit scores $S$ produced by PET’s probing methods (i.e., Algorithms 2 and 1) is order-preserving (isotonic) with respect to the true latent attractiveness scores $\bm{q}$. Formally, for any two clusters $c_{i},c_{j}$, we have $q_{i}\geq q_{j}\Rightarrow\mathbb{E}[S[i]]\geq\mathbb{E}[S[j]].$
1 formalizes the goal that a successfully trained model should learn scores that are correctly ordered according to the true underlying preferences, which is a foundation for many modern preference learning models (guo2017calibration; desai2020calibration; kadavath2022language; schick2020exploiting). With this, we can state the following:
Lemma 1 (Optimality of PET’s Ranking).
With 1, the ranking produced by PET by sorting its inferred probabilities maximizes the expected score for common order-aware metrics, including NDCG@k, Recall@k, and Precision@k. (See proof at Appendix D.)
Remark 1 (Sub-optimality of Direct Generation).
The sub-optimality of direct generation stems from its autoregressive nature, where the ranking is an emergent property of a sequence generation task rather than a direct readout of preference scores. This process is susceptible to multiple known issues: exposure bias, where the model’s training distribution differs from its inference distribution (ranzato2015sequence); calibration errors, where output probabilities are skewed by factors like token frequency (zhao2021calibrate); and search errors from the decoding heuristics themselves, such as beam search (wiseman2016sequence).
4 Experiment
4.1 Experiment Design and Setup
Our experimental design is crafted to comprehensively evaluate the efficacy of PET’s ability in discerning and predicting user preferences across varying granularities and long-term and short-term dynamics. We focus on using two primary inference methodologies for preference prediction and benchmark them against established and SOTA baselines.
Datasets. We evaluate PET on three real-world datasets of varying scale and complexity. MovieLens (harper2015movielens): a widely-used dataset for movie recommendations, representing a relatively small and well-defined genre (cluster) space with 19 clusters (i.e., movie genres). In total, we used 200k users with 33 million rating histories associated with 87k movies. Yelp: a large-scale dataset with 2 million users and 8 million ratings across 1,311 business categories, providing a significantly higher-dimensional preference space. A Real-World Short-Video Dataset: this proprietary dataset is derived from the a real-world short video platform and captures large-scale real-world user interaction behaviors in a highly dynamic environment, reflecting diverse and rich user preference patterns. To ensure privacy, all identifiers are rigorously anonymized. The dataset comprises 50 million implicit feedback records from 16k randomly sampled users, covering 15 million items. Each record contains only anonymized user IDs, item IDs, and interaction types.
[…]
Models and Methods. We use two families of open-source 8B-parameter language models: Qwen3-8B (qwen3technicalreport) and DeepSeek-Distill-Llama-8B (deepseekai2025deepseekr1incentivizingreasoningcapability), selected for their strong balance of SOTA performance and computational efficiency. Each model is trained with our PT+SFT (LoRA fine-tune (hu2022lora)) pipeline. This setup supports a comprehensive ablation of training strategies. Our experiments evaluate the primary inference methodologies detailed in Section 3: Likelihood-based Probing (Algorithm 1) and Generative Classification (Algorithm 2). For the Yelp dataset with large cluster space, we specifically evaluate our Hierarchical Probing approach (Algorithm 3). For all methods, the final inferred probability distribution is used to generate a ranked list of clusters for evaluation.
Baselines. We compare PET with two strong LLM-based baselines. To isolate the effect of our inference paradigm, the first baseline Direct Generation uses the same pre-trained and fine-tuned LLM as PET but is prompted to directly generate a top-k ranked list of preferences without producing a probability distribution (geng2022recommendation; deng2025onerec; wang2025user; wang2024llms). As this method does not produce a probability distribution, it is evaluated only on ranking metrics. The second baseline is Qwen3-Reranker-8B, a SOTA pre-trained model specifically optimized for ranking tasks (qwen3embedding). We use it out of the box without any additional fine-tuning. Given a user history and a candidate cluster space, it produces both a ranked list and a probability distribution. For the short-video dataset, we compare against the platform’s production Transformer-based ranker model.
Evaluation Metrics. Model performance is comprehensively evaluated using established metrics compared to long-term (LT) and short-term (ST) ground truths, with a focus on ranking quality and distributional similarity. Ranking metrics include NDCG@k, which assesses the quality of ranking order with an emphasis on top positions; Recall@k, measuring the proportion of relevant items captured within the top k results; and Precision@k, evaluating the exactness of top-k recommendations. Distribution similarity is quantified using Jensen-Shannon Divergence (JS-Div), where lower values indicate closer alignment between predicted and actual preference distributions. All results are averaged across test samples and reported per context window size as specified in the dataset.
Figure 2: Comprehensive comparison of different methods on MovieLens dataset across a range of context windows (i.e., 1,3,5, and 8 sessions). Prediction windows: long-term and short-term. Metrics: NDCG@$1,5,10$ and JS-Divergence. We compare our Logit Probing and Generative Classification methods against (1) Direct Generation benchmark trained (PT+SFT) on a Qwen3-8B base model and (2) a SOTA Qwen3-Reranker-8B ranking model. Note that for the Direct Generation baseline, the JS-divergence will not be available. Sample size: 38,434.
4.2 How Effective and Robust is PET across Different Context Horizon?
We begin by evaluating PET on the MovieLens dataset with 19 clusters, aiming to systematically test its effectiveness and robustness across following key dimensions: (1) different input history lengths and (2) prediction horizons (short-term vs. long-term). We vary the input context window from 1 to 8 sessions, and assess performance on both short-term (next-session) and long-term (multi-session) preference prediction.
As shown in Figure 2, the results demonstrate the robustness and effectiveness of PET across a wide range of settings. The performance gains are especially pronounced in full-list metrics (NDCG@5, NDCG@10), where PET’s distributional inference enables richer modeling of user preferences beyond top-1 accuracy. For example, with Qwen3-8B, context window =8=8, and long-term prediction, Logit-Probing achieves NDCG@10 = 0.863, with a +55% improvement over Direct Generation and +20% gain over Qwen3-Reranker-8B. It also improves distributional alignment, reducing JS-Div to 0.316, a 22% relative reduction compared to the Qwen3-Reranker-8B’s 0.406. While Direct Generation is occasionally competitive at NDCG@1, its performance deteriorates as the number of predicted clusters increases, as it might suffers from exposure bias, noted in Remark 1. Our Generative Classification method also shows consistent improvements over Direct Generation on Qwen3-8B base model. Though slightly less accurate than Logit-Probing, it offers a compelling latency–accuracy trade-off due to its single-pass nature, making it a strong candidate for real-time or resource-constrained environments. (See more experimental results in other settings in Appendix F)
To demonstrate PET’s ability to build interpretable profiles, we sampled 125 users and divided their histories into four equal time periods. Using PET’s Likelihood-based Probing, we inferred a distribution over 19 genres for each user in each period. Aggregating these across users yielded group-level preference distributions that evolve over time – visualized in both a heatmap and 3D ribbon plot (Figure 3). These results show that user preference distributions are not static but migrate in nuanced ways. Dominant genres like Drama, Comedy, and Action exhibit stability across periods, representing structural, long-term user interests. In contrast, genres such as Adventure, Romance, and Sci-Fi demonstrate significant temporal fluctuations – highlighting PET’s ability to capture volatile, context-dependent preferences. Notably, genres like Sci-Fi, Fantasy, and IMAX show gradual upward shifts, possibly reflecting emerging or re-surfacing interests. Together, these dynamics validate PET’s unique capability not only in preference prediction, but also in generating descriptive, evolving group-level user profiles – an essential feature for diagnosing user behavior shifts, designing cold-start interventions, or understanding preference stability versus exploration.
4.3 Can PET Handle Thousands of Clusters?
To evaluate PET’s scalability in high-dimensional settings, we use the Yelp dataset with 1,311 fine-grained categories. Since exhaustive Logit-Probing becomes intractable at this scale, we adopt a two-stage Hierarchical Probing strategy (Algorithm 3). First, we construct a semantic hierarchy that maps all categories into 26 high-level L1 clusters (e.g., “Food & Restaurants,” “Health & Medical”). For demonstration, we focus on the high-traffic L1 cluster “Food & Restaurants,” further decomposed into 19 L2 sub-clusters (e.g., “Mexican,” “Vegan”). During inference, PET first identifies a user’s coarse L1 interest, then performs fine-grained ranking within the selected L2 cluster. While operators may target different L1 clusters depending on downstream needs, this experiment provides one representative use case. (See detailed mapping construction in Appendix E.)
Table 1 shows that PET with Hierarchical Probing achieves strong performance across both coarse (L1) and fine-grained (L2) levels. At L1, Logit-Probing is unequivocally dominant. It achieves high and robust NDCG scores (e.g., NDCG@1,5,10 are 0.980, 0.879, and 0.915 for Qwen3-8B), demonstrating an exceptional ability to identify a user’s general interests. In stark contrast, both the Direct Generation and Generative Classification methods fail at this stage, with performance often near zero. These results reinforce the necessity of iterative probing in high-dimensional settings where broad disambiguation is critical. At L2, where the ranking task becomes more localized, both Logit-Probing and Generative Classification significantly outperform the Direct Generation benchmark. Notably, while Logit-Probing remains the top performer for Qwen3-8B (0.666 NDCG@10), Generative Classification emerges as the clear winner for the DeepSeek-distill-Llama-8B model (0.574 NDCG@10), suggesting that decoding strategies for fine-grained inference can be model-dependent.
Together, these findings demonstrate that PET, equipped with hierarchical probing, scales effectively to large cluster spaces. The coarse-to-fine process not only preserves tractability but also allows flexible adaptation across model architectures — with Logit-Probing excelling in breadth and Generative Classification offering a strong, efficient alternative at finer levels of granularity.
(a) Group-level preference evolution as a probability heatmap
(b) Group-level preference evolution as a bar chart
Figure 3: Evolution of group-level movie genre preferences as probability distribution across four time periods. Dataset: Movielens, 125 users over 4 periods. Method: PET Likelihood-based Probing with PT+SFT on Qwen3-8B. Note: due to limited space, we only show top-10 clusters (genres) here.
Qwen3-8B DeepSeek-distill-Llama-8B
Level Method NDCG@1 NDCG@5 NDCG@10 NDCG@1 NDCG@5 NDCG@10
L1 (26 clusters) Direct Generation .008 .139 .330 .016 .308 .343
Logit Probing .980 .879 .915 .925 .856 .880
Generative Classification .107 .118 .163 .001 .062 .119
L2 (19 sub-clusters) Direct Generation .298 .368 .509 .200 .153 .329
Logit Probing .513 .562 .666 .261 .268 .363
Generative Classification .466 .349 .454 .684 .553 .574
Table 1: Hierarchical Probing’s performance on the high-dimensional Yelp dataset for the long-term prediction task. Results are shown for a context window of 8 using the PT+SFT pipeline. Best NDCG scores for each level are in bold. L1 category contains 26 clusters; L2 category contains 19 clusters under one L1 category of “Food & Restaurants”. Alternative taxonomies can be easily integrated depending on platform objectives. Sample Size: 5000.
4.4 How Does PET Perform on Real-World Short-Video Platform?
We conclude our empirical study with a large-scale, private short-video dataset from a major content platform, comprising 78 content clusters. This experiment is designed to test our method’s efficacy in a real-world environment against a heavily optimized, production-level Transformer-based ranking model. The primary challenge in such an ecosystem is overcoming the inherent feedback loop, where users primarily interact with content they are already shown. Therefore, our evaluation focuses on PET’s ability to complement an existing production model that is heavily optimized for popular, high-frequency head content. In contrast, PET is designed to surface niche, long-tail interests that are often suppressed by popularity bias. We aggregate interactions into daily sessions (a 24-hour window), ensuring a denser and more comprehensive view of a user’s preferences. All results use the PT+SFT pipeline with Qwen3-8B. Following industry practice, we use play duration – the total time a user spends watching each content item – as a robust and continuous proxy for engagement. (See more details in Section E.2)
Results in Table 2 showcase PET’s effectiveness in promoting long-tail interest, diversity, and personalized recommendation. We highlight the setting of 30-day history to 14-day prediction, as it represents the most challenging long-range forecasting task. The production model fails to identify users’ long-tail preferences, with an NDCG@20 of just 0.0243 on the long-tail segment. PET with Logit-Probing achieves NDCG@20 of 0.1971 on the same tail – a +711% improvement. This demonstrates PET’s ability to surface underrepresented, user-specific content that is typically overlooked by popularity-driven production models. Notably, results are consistent for $k\in\{1,5,10,20\}$ with PET achieving more gains compared to the production model when $k$ is small. We also compare across prediction horizons and observe that PET slightly outperforms the baseline more significantly when the forecasting window is longer (14 days vs. 7 days), consistent with our earlier findings at Section 4.2 regarding LLMs’ strength in modeling long-term user preferences. Finally, PET’s performance remains robust across different context lengths. It maintains stable ranking quality even with limited user history, underscoring its ability to effectively summarize user interests with sparse or noisy data.
Taken together, these findings validate PET as a practical and scalable complement to production systems. enhancing long-tail coverage and personal relevance without sacrificing overall performance. Its ability to enhance diversity and personalization – especially for the most engaged users – marks a valuable addition to real-world recommender systems.
Context Prediction Method NDCG@1 NDCG@5 NDCG@10 NDCG@20
14 Days 7 Days SOTA Production (Production) .007 .009 .012 .034
Logit-Probing .113 .124 .147 .203
14 Days SOTA Transformer (Production) .004 .006 .008 .024
Logit-Probing .110 .126 .150 .208
30 Days 7 Days SOTA Transformer (Production) .007 .009 .012 .034
Logit-Probing .103 .118 .138 .192
14 Days SOTA Transformer (Production) .004 .006 .008 .024
Logit-Probing .108 .122 .143 .197
Table 2: Long-tail preference learning performance on the short-video dataset for high-activity users, comparing PET’s Logit Probing against a SOTA production model across various history (i.e., 14 and 30 days) and prediction windows (i.e., 7 and 14 days). All results are from the PT+SFT pipeline with the Qwen3-8B model.
@@ 5 Conclusion
In this paper, we introduce Preference Evolution Tracking (PET), a framework that casts LLM-based personalization as distributional preference mapping: PET infers a user’s cluster-level probability vector via Likelihood-based Probing and Generative Classification, and scales to large label spaces through Hierarchical Probing. Under a mild isotonicity assumption, ranking by PET’s inferred probabilities is optimal for standard order-aware metrics. Empirically, PET is robust across datasets and horizons, outperforming strong generative and reranking baselines and substantially improving coverage of long-tail interests on an industrial-scale corpus. Beyond accuracy, PET yields interpretable, temporally evolving user profiles that enable group-level user analysis and preference-shift detection. Future work includes extending PET to model richer user representations by incorporating auxiliary signals such as demographic attributes, behavioral embeddings, and multimodal content. These enhancements aim to enable more personalized, adaptive, and interpretable preference modeling in real-world recommendation systems.
@@ Appendix A Algorithms
16:  $S[j]\leftarrow\frac{\exp(s_{+})}{\exp(s_{+})+\exp(s_{-})}$ $\triangleright$ Compute ’yes’ probability as the cluster’s score
@@
7:Predicted preferences $\hat{\theta}$ over the L2 clusters $C_{L2}$.
@@
10:Use $H_{1:t}$ and $C_{L1}$ to get L1 scores $S_{L1}$ via “Likelihood-based Probing” (Algorithm 1).
@@ Appendix B Prompt of Inference
Prompt for Likelihood-based Probing Algorithm 1
@@
Prompt for Generative Classification Algorithm 2
@@ Appendix C Computational Complexity Analysis
Let $K$ be the number of preference clusters and $T$ the prompt length (tokens). Let $C(\mathcal{M},T)$ denote the cost of a single forward pass of the language model $\mathcal{M}$ on a prompt of length $T$ (the dominant compute). Each cluster $c$ is associated with a fixed “verbalizer” (token or short phrase) used to read its score from next-token logits.
Complexity of PET Methods
Likelihood-based Probing Algorithm 1.
Our implementation is iterative: we issue a focused prompt per cluster, effectively scoring clusters independently. The total cost scales linearly with the number of probes:
@@
Generative Classification Algorithm 1.
@@
Hierarchical Probing Algorithm 3.
@@
$\text{Cost}\;\approx\;O\!\big(K_{1}\cdot C(\mathcal{M},T_{L1})\big)\;+\;O\!\big(B\cdot K_{2,\text{avg}}\cdot C(\mathcal{M},T_{L2})\big).$
@@
One prompt with autoregressive decoding of approximately $O(k)$ output tokens:
$\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T)+k\big).$
If re-prompted separately for multiple cutoffs (e.g., $k=1,5,10$), end-to-end latency increases linearly with the number of cutoffs.
@@
  • Multi-$k$ evaluation. PET’s distributional methods (Generative Classification; Hierarchical with small $B$) yield a full distribution in the same pass(es), so reporting multiple $k$ does not require re-decoding.
Appendix D Proofs
For Lemma 1:

### v3

@@ 5 Conclusion
We presented Preference Evolution Tracking (SPECTRA), a framework for distributional user modeling with LLMs. Given a user’s history, SPECTRA infers an explicit posterior over interpretable preference clusters via Likelihood-based Probing or Generative Classification, and scales to large cluster spaces with Hierarchical Probing. In theory, SPECTRA decouples preference inference from sequence decoding by ranking clusters using explicit probability readouts, which avoids decoding-induced exposure bias that can degrade Direct Generation at larger cutoffs. Across public benchmarks and an industrial short-video dataset, SPECTRA is robust across context lengths and time scales, consistently improving distributional alignment, mitigating popularity concentration, and strengthening cluster-level ranking compared to strong baselines. Beyond predictive gains, SPECTRA produces inspectable, temporally evolving preference profiles that support auditing and group-level analysis of preference shifts. Future work includes extending the framework to incorporate additional signals (e.g., multimodal content and auxiliary behavioral representations) and studying robustness under evolving taxonomies and feedback loops in deployed personalization pipelines.
@@ Appendix A Algorithms
16:  $S[j]\leftarrow s_{+}-s_{-}$ $\triangleright$ Compute ’yes’ probability as the cluster’s score
@@
7:Predicted preferences $\hat{\theta}_{t+1:T}$ over the L2 clusters $C_{L2}$.
@@
10:Use $H_{1:t}$ and $C_{L1}$ to get L1 scores $S_{L1}$ via “Likelihood-based Probing” (Algorithm˜1).
@@ Appendix B Prompt of Inference
Prompt for Likelihood-based Probing Algorithm˜1
@@
Prompt for Generative Classification Algorithm˜2
@@ Appendix C Computational Complexity Analysis
Let $K$ be the number of preference clusters and $T$ the total prompt length (tokens). Let $C(\mathcal{M},T)$ denote the cost of a single forward evaluation of the language model $\mathcal{M}$ on $T$ tokens (dominant compute). For KV/prefix reuse, we decompose each probing prompt into a shared prefix of length $T_{\mathrm{pre}}$ (user history + fixed scaffold) and a cluster-specific suffix of length $T_{\mathrm{suf}}$, so that $T=T_{\mathrm{pre}}+T_{\mathrm{suf}}$. We write $C(\mathcal{M},T_{\mathrm{pre}})$ and $C(\mathcal{M},T_{\mathrm{suf}})$ for the corresponding costs.
Verbalizers. In generative classification, each cluster $c$ is associated with a verbalizer (e.g., a single token/letter) whose next-token logit is read and normalized. In likelihood probing, the cluster is inserted into the prompt and we read the next-token logits of a fixed response verbalizer set (e.g., Yes/No).
Complexity of SPECTRA Methods
Likelihood-based Probing Algorithm˜1.
Our probing implementation is iterative: we score clusters via $K$ cluster-conditioned prompts. Without reuse, this requires $K$ full prompt evaluations:
@@
With prefix/KV reuse, we evaluate the shared prefix once and only re-evaluate a short cluster-specific suffix per probe:
$\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T_{\mathrm{pre}})+K\cdot C(\mathcal{M},T_{\mathrm{suf}})\big).$
[…]
Generative Classification Algorithm˜2.
@@
Hierarchical Probing Algorithm˜3.
@@
Cost $\;\approx\;O\!\big(K_{1}\cdot C(\mathcal{M},T_{L1})\big)$
$\quad\;+\;O\!\big(B\cdot K_{2,\text{avg}}\cdot C(\mathcal{M},T_{L2})\big).$
@@
One prompt with an initial prefill plus autoregressive decoding of $m(k)$ output tokens (typically $m(k)=O(k)$ under a constrained letter-list format):
$\text{Cost}\;\approx\;O\!\big(C(\mathcal{M},T)+m(k)\cdot C_{\mathrm{dec}}(\mathcal{M})\big),$
where $C_{\mathrm{dec}}(\mathcal{M})$ denotes the per-token decode cost given a cached prefix (implementation-dependent). If re-prompted separately for multiple cutoffs (e.g., $k=1,5,10$), end-to-end cost scales linearly with the number of cutoffs.
@@
Practical overhead (token-length proxy).
Define the relative compute overhead of probing (with reuse) versus a single-shot method as
$\rho\;=\;\frac{C(\mathcal{M},T_{\mathrm{pre}})+K\,C(\mathcal{M},T_{\mathrm{suf}})}{C(\mathcal{M},T_{\mathrm{pre}}+T_{\mathrm{suf}})}.$
In typical prompting setups, $T_{\mathrm{pre}}$ (user history) can dominate and $T_{\mathrm{suf}}$ is relatively short, so prefix reuse can reduce the overhead from $\approx K\times$ to a constant factor that depends on the ratio $T_{\mathrm{pre}}/T_{\mathrm{suf}}$. Absolute latency depends on batching and serving implementations; the proxy above is intended to summarize scaling rather than provide measured runtimes.
Method # model evals No reuse With prefix/KV reuse
Direct Generation (top-$k$) 11 prefill + $m(k)$ decode $\approx O(C(\mathcal{M},T)+m(k)\cdot C_{\mathrm{dec}}(\mathcal{M}))$ (same)
Generative Classification 11 $O(C(\mathcal{M},T))$ (same)
Likelihood Probing $K$ $O(K\,C(\mathcal{M},T))$ $O(C(\mathcal{M},T_{\mathrm{pre}})+K\,C(\mathcal{M},T_{\mathrm{suf}}))$
Table 4: Complexity summary. $m(k)$ is the number of generated tokens in direct generation.
Example regimes (token-length proxy).
For illustration, consider K=20K{=}20 and compare probing with reuse ($T_{\mathrm{pre}}+KT_{\mathrm{suf}}$) to a single-shot pass ($T_{\mathrm{pre}}+T_{\mathrm{suf}}$) under two regimes: (i) short history: $T_{\mathrm{pre}}{=}80$, $T_{\mathrm{suf}}{=}20$, giving a ratio $\frac{80+20\cdot 20}{80+20}=\frac{480}{100}\approx 4.8\times$; (ii) long history: $T_{\mathrm{pre}}{=}300$, $T_{\mathrm{suf}}{=}15$, giving $\frac{300+20\cdot 15}{300+15}=\frac{600}{315}\approx 1.9\times$. These token-length ratios are proxies for scaling; actual wall-clock depends on batching and kernels.
Regime $T_{\mathrm{pre}}$ $T_{\mathrm{suf}}$ $K$ Proxy ratio $\frac{T_{\mathrm{pre}}+KT_{\mathrm{suf}}}{T_{\mathrm{pre}}+T_{\mathrm{suf}}}$
[…]
Table 5: Token-length proxy ratios for probing with prefix/KV reuse (illustrative; not measured latency).
@@
  • Multi-$k$ evaluation. SPECTRA’s distributional methods (Generative Classification; Hierarchical with small $B$) yield a full distribution in the same pass(es), so reporting multiple $k$ does not require re-decoding.
Appendix D Extra Theories and Proofs
For the theoretical analysis, we consider an idealized preference mapping $g_{\phi}:H_{1:t}\rightarrow\mathbb{R}^{K}$ that produces logits $\boldsymbol{S}(H_{1:t})=g_{\phi}(H_{1:t})$ and an induced distribution $\hat{\theta}_{t+1:T}(H_{1:t})=\mathrm{softmax}(\boldsymbol{S}(H_{1:t}))$. We assume that $\phi$ is chosen to minimize the population cross-entropy between the latent-softmax model and the predictor:
$\mathcal{L}_{\mathrm{CE}}(\phi)=\mathbb{E}_{H_{1:t}}\big[\mathrm{CE}\big(\theta_{t+1:T},\hat{\theta}_{t+1:T}\big)\big].$ (2)
Training on empirical labels $\bar{\theta}_{t+1:T}$ can be viewed as a finite-sample approximation to this population objective. For the theoretical analysis in this appendix, we work in an idealized population setting. We posit a latent preference distribution $\theta(H_{1:t})=\mathrm{softmax}(\mathbf{q}(H_{1:t}))$ and view $\bar{\theta}_{t+1:T}$ as a finite-sample Monte Carlo estimate of $\theta(H_{1:t})$ (i.e., $\bar{\theta}_{t+1:T}\approx\theta(H_{1:t})$ as the number of interactions grows). Consequently, the empirical objective with $\bar{\theta}$ approximates the population objective with $\theta$.
D.1 Preliminaries: Idealized Bayes-optimality
We first establish the behavior of the model in an idealized population limit. Let $\boldsymbol{S}\in\mathbb{R}^{K}$ be the model’s logits, produced by SPECTRA’s probing methods (i.e., Algorithms˜2 and 1). The model predicts $\hat{\theta}=\mathrm{softmax}(\mathbf{S})$, and we train by minimizing cross-entropy between $\theta$ and $\hat{\theta}$. This follows the standard view of cross-entropy (log-loss) as a strictly proper scoring rule: at the population optimum, the predictor’s output distribution coincides with the true conditional distribution (Gneiting and Raftery, 2007; Goodfellow et al., 2016; Blasiok et al., 2023). In multiclass settings, this implies Bayes-consistency of the induced classifier (Tewari and Bartlett, 2007).
Proposition 1 (Isotonicity at convergence).
Since cross-entropy is a strictly proper scoring rule (Gneiting and Raftery, 2007; Blasiok et al., 2023), the global minimum of the population risk is achieved if and only if $\mathrm{softmax}(\boldsymbol{S})=\mathrm{softmax}(\boldsymbol{q})$. Equivalently, in the latent-softmax model $\theta=\mathrm{softmax}(\boldsymbol{q})$, the Bayes-optimal predictor recovers the true conditional distribution (Tewari and Bartlett, 2007; Goodfellow et al., 2016). By injectivity of softmax up to an additive constant, this implies $\boldsymbol{S}=\boldsymbol{q}+c\boldsymbol{1}$ for some scalar $c$. Thus, in the idealized limit, the model’s logits are perfectly order-preserving (isotonic) with respect to latent preferences.